# Field of values analysis that includes zero for preconditioned nonsymmetric saddle-point systems

### v1

@@ Abstract
We present a field-of-values (FOV) analysis for preconditioned nonsymmetric saddle-point linear systems, where zero is included in the field of values of the matrix. We rely on recent results of Crouzeix and Greenbaum [Spectral sets: numerical range and beyond. SIAM Journal on Matrix Analysis and Applications, 40(3):1087-1001, 2019], showing that a convex region with a circular hole is a spectral set. Sufficient conditions are derived for convergence independent of the matrix dimensions. We apply our results to preconditioned nonsymmetric saddle-point systems, and show their applicability to families of block preconditioners that have not been previously covered by existing FOV analysis. A limitation of our theory is that the preconditioned matrix is required to have a small skew-symmetric part in norm. Consequently, our analysis may not be applicable, for example, to fluid flow problems characterized by a small viscosity coefficient. Some numerical results illustrate our findings.
Keywords. field of values, nonsymmetric saddle-point systems, GMRES convergence, block preconditioners
@@ 1 Introduction
$W(A)=\left\{\frac{x^{*}Ax}{x^{*}x}:\quad x\in\mathbb{C}^{n}\right\}$
and the $H$-field of values of $A$, given another matrix $H\in\mathbb{C}^{n\times n}$, is defined as
$W_{H}(A)=\left\{\frac{x^{*}HAx}{x^{*}Hx}:\quad x\in\mathbb{C}^{n}\right\}.$
Early work on the topic was published in [11, 16] and in several other papers; see [3] for a recent expository paper that provides an overview of the use of FOV, its history and development, and a comprehensive list of references.
$\begin{bmatrix}F&B^{T}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ p\end{bmatrix}=\begin{bmatrix}f\\ g\end{bmatrix},$ (1)
where $F\in\mathbb{R}^{n\times n}$ is nonsymmetric, $B\in\mathbb{R}^{m\times n}$ has full row rank, and $u,f\in{\mathbb{R}}^{n},\ p,g\in{\mathbb{R}}^{m}.$
Significant work has been done on this topic [2, 4, 16, 17, 19], but to the best of our knowledge, the analysis is limited to the situation where 00 is not included in the field of values. Our goal is to perform an FOV analysis for preconditioned saddle-point systems in the case where the origin is included.
Part of our motivation in considering the field of values is that spectral analysis may be limited for this family of linear systems: for nonsymmetric saddle-point systems arising from PDEs, the condition number of the eigenvector matrix of the preconditioned matrix typically increases with the matrix dimensions. In the context of the Navier-Stokes equations, for example, this happens even for a large viscosity coefficient. Thus, an analysis of the eigenvalues of the preconditioned matrix is often insufficient to theoretically prove convergence for nonsymmetric saddle-point matrices.
@@
Given a Hermitian positive definite matrix $H\in\mathbb{C}^{n\times n}$, the $H$-norm of a vector $u\in\mathbb{C}^{n}$ is defined as
$\|u\|_{H}=(u,Hu)^{1/2}.$
@@
[7] Let $r_{k}=b-Ax_{k}$ be the residual of the $k$-th iteration, $x_{k}$, of GMRES with respect to the $H$-norm applied to the linear system $Ax=b$ of the residual, and let $\mathbf{P}_{j}$ denote all polynomials $p$ of degree $\leq j$ that satisfy $p(0)=1$. Then,
$\frac{\|r_{j}\|_{H}}{\|r_{0}\|_{H}}\leq\min_{p\in\mathbf{P}_{j}}\|p(A)\|_{H}\leq(1+\sqrt{2})\min_{p\in\mathbf{P}_{j}}\max_{z\in W_{H}(A)}|p(z)|.$
A challenge is that when $0\in W_{H}(A)$, we have $\min_{p\in\mathbf{P}_{j}}\max_{z\in W_{H}(A)}|p(z)|=1$, and Theorem 1.4 fails to provide a useful bound on GMRES convergence in this case. The presence of a zero in the field of values is, in fact, common in saddle-point systems: the (2,2)-block of a saddle-point system preconditioned with a block-diagonal matrix can be 00. Recently, Crouzeix and Greenbaum [7] have defined a convex region with a circular hole and showed that it is a spectral set. This can be used to analyze cases when zero is included in the field of values.
@@
[7] Let $A$ be a matrix of the same dimensions as $H$. Then, $\Omega_{CG}=W_{H}(A)\cap\{z\in C:|z|\geq\|A^{-1}\|_{H}^{-1}\}$ is a $(3+2\sqrt{3})$-spectral set for $A$.
@@
The field of values of a matrix is difficult to compute, and in the context of the iterative solution of linear systems, it is necessary to exploit the specific properties of the matrices involved in order to provide concrete conditions for scalability of an iterative solver. Our work extends the family of saddle-point linear systems for which FOV analysis is applicable. In particular, we consider block-diagonal preconditioners and certain block-triangular preconditioners for which no previous FOV analysis is available. On the other hand, our analysis has some limitations compared to the well-studied FOV analysis that excludes the origin. For example, in [17], scaling is effectively used to allow for applying FOV analysis to the discrete Navier-Stokes equations with a small viscosity coefficient when the field of values does not include the origin. In our analysis we are not able to utilize scalings in the same manner, and we require the skew-symmetric part of the linear system to be small norm-wise.
The remainder of this paper is structured as follows. In Section 2 we present an analysis that deals with zero in the field of values. In Section 3 we specialize our results to saddle-point systems. In Section 4 we discuss a few examples of relevant applications and present some numerical results. Finally, we draw some conclusions in Section 5.
@@ 2.1 Preliminaries
Let us present a few known results that we will use in our analysis.
@@
For two symmetric positive definite matrices $H_{1}\in\mathbb{R}^{n\times n}$ and $H_{2}\in\mathbb{R}^{m\times m}$, we define the $(H_{1},H_{2})$-norm for a matrix $M\in\mathbb{R}^{m\times n}$ as
$\|M\|_{H_{1},H_{2}}=\max_{v\neq 0}\frac{\|Mv\|_{H_{2}}}{\|v\|_{H_{1}}}.$
It is immediate from Definition 2.1 that
@@
The following properties from [17] are useful for our analysis.
Lemma 2.2 ([17, Lemma 1]).
@@
Let $P_{n}$ denote the set of polynomials $p$ of degree at most $n$ with $p(0)=1$. For a compact set $S$ in the complex plane, with the origin not included in or surrounded by $S$ and no isolated points, define
$E_{n}(S)=\min_{p\in P_{n}}\max_{z\in S}|p(z)|$
@@
Let $g(z)$ be the Green’s function associated with $S$, defined in the exterior of $S$, satisfying $\nabla^{2}g=0$ outside of $S$, $g(z)\to 0$ as $z\to\partial S$, and $g(z)-\log|z|\to C$ as $|z|\to\infty$ for some constants $C$. Then,
@@ 2.2 Sufficient Conditions
Let $A,H\in\mathbb{R}^{n\times n}$ where $A$ is nonsingular and $H$ is symmetric positive definite. Then, GMRES converges with respect to $H$-norm in a fixed number of iterations independent of its dimension, $n$, if the following conditions hold for some constants $a$, $b$, and $c$:
@@
$\|(HA-A^{T}H)/2\|_{H,H^{-1}}\leq c;$ (2c)
$bc<1.$ (2d)
@@
$|Im(z)|$ $\leq\max_{x\in\mathbb{C}^{n}}\left|\left(\frac{x^{*}HAx}{x^{*}Hx}-\left(\frac{x^{*}HAx}{x^{*}Hx}\right)^{*}\right)\Big{/}2\right|$
$=\max_{x\in\mathbb{C}^{n}}\left|\left(\frac{x^{*}(HA-A^{T}H)x}{2x^{*}Hx}\right)\right|$
@@
Then,
@@
$\frac{\|r_{j}\|_{H}}{\|r_{0}\|_{H}}$ $\leq\min_{p\in\mathbf{P}_{j}}\|p(A)\|_{H}\leq(3+2\sqrt{3})\min_{p\in\mathbf{P}_{j}}\max_{z\in\Omega_{CG}}|p(z)|.$
$\text{Im}(z)=c$$\text{Im}(z)=-c$$\text{Re}(z)$$\text{Im}(z)$
Figure 1: The shaded region is $\Omega_{D}$ when conditions (2a)–(2d) of Lemma 2.4 hold
Since Condition (2d) holds, the origin is not surrounded by $\Omega_{CG}$, and it follows from Theorem 2.3 that there is always a polynomial of (of some degree) with value 1 at the origin that has a maximum magnitude strictly less than 1 on the closure of this set and hence GMRES converges with an asymptotic rate given by $\exp(-g(0))<1$, where $g$ is the Green’s function of this set with a pole at $\infty$ [6, 7]. ∎
@@
If condition (2d) of Lemma 2.4 does not hold, the iterative solver may still converge but we cannot prove convergence using our technique of proof. Specifically, it is immediate to see that $\Omega_{CG}$ is connected and due to the maximum modulus principle, we can only obtain $\min_{p\in\mathbf{P}_{j},p(0)=1}\max_{z\in\Omega_{CG}}|p(z)|\geq 1$, which does not indicate convergence; see Figure 2 for a graphical illustration.
$\text{Im}(z)=c$$\text{Im}(z)=-c$$\text{Re}(z)$$\text{Im}(z)$
Figure 2: The shaded region is $\Omega_{D}$ when $bc\geq 1$ (i.e., when condition (2d) of Lemma 2.4 is violated)
@@ 2.3 Scope and Limitations of the Analysis
Recall a widely used definition of FOV equivalence (see, for example, [17]).
@@
Given two nonsingular matrices $M,A\in\mathbb{R}^{n\times n}$, $M$ is $H$-field-of-values equivalent to $A$ if there exist positive constants $\alpha,\beta$ independent of $n$ such that
$\alpha\leq\frac{(MAx,x)_{H}}{(x,x)_{H}},\ \frac{\|MAx\|_{H}}{\|x\|_{H}}\leq\beta.$ (3)
@@
In contrast, in our case, condition (2d) requires $bc$ to be small. While scaling reduces one of $b$ or $c$, it increases the other. Therefore, a simple scaling strategy does not work in the case we are considering, which reveals a limitation of our analysis. We note that condition (2b) is rather standard by norm equivalence considerations (see, for example, [17, Lemma 3]). It is condition (2c) that seems to present the difficulty, because it requires the skew-symmetric part of the operator to be smaller than the radius of the inner disk; see Figure 1. Therefore, practically speaking, our analysis is limited to cases where the preconditioned matrix is only mildly nonsymmetric.
However, we note that this lemma can be improved to allow for looser conditions by using a more sophisticated analysis.
@@
The field of values of $A$ is a convex hull of two disks centered at $-1$ with radius $1/4$ and at $2$ with radius $1.2$, independently of the matrix dimensions. The inverse $A^{-1}$ is available analytically, and it can be shown that $\|A^{-1}\|_{2}^{-1}\to\frac{3}{4}$ as $n\to\infty$; see, for example, [15] for useful relevant results for Toeplitz matrices. For a finite value of $n$, the norm needs to be computed numerically, and we have experimentally observed that it is bounded between $0.74$ and $0.76$ for relatively modest values of $n$.
We provide a graphical illustration in Figure 4. Here $c=1.2$ and $b\geq\frac{1}{0.76}$. The condition (2d) is violated, but GMRES would still converge for a linear system with the matrix $A$ because $\Omega_{CG}$ does not surround/include the origin. A more careful analysis that tracks the boundary of the FOV (see, e.g., [18]) might result in conditions that are easier to satisfy.
@@
While the limitation we have noted is considerable, our analysis substantially broadens the scope of preconditioners for which FOV analysis can be carried out. In particular, in terms of the quantities of Definition 2.5, our analysis makes it possible to consider
$\frac{(MAx,x)_{H}}{(x,x)_{H}}\leq 0.$ (4)

### v2

# GMRES Convergence Analysis for Nonsymmetric Saddle-Point Systems When the Field of Values Contains the Origin

@@ Abstract
We present a field-of-values (FOV) analysis for preconditioned nonsymmetric saddle-point linear systems, where zero is included in the field of values of the matrix. We rely on recent results of Crouzeix and Greenbaum [Spectral sets: numerical range and beyond. SIAM Journal on Matrix Analysis and Applications, 40(3):1087-1101, 2019], showing that a convex region with a circular hole is a spectral set. Sufficient conditions are derived for convergence independent of the matrix dimensions. We apply our results to preconditioned nonsymmetric saddle-point systems, and show their applicability to families of block preconditioners that have not been previously covered by existing FOV analysis. A limitation of our theory is that the preconditioned matrix is required to have a small skew-symmetric part in norm. Consequently, our analysis may not be applicable, for example, to fluid flow problems characterized by a small viscosity coefficient. Some numerical results illustrate our findings.
Keywords. field of values; nonsymmetric saddle-point systems; GMRES convergence; block preconditioner
@@ 1 Introduction
[…]
and the $H$-field of values of $A$, given a Hermitian positive definite matrix $H\in\mathbb{C}^{n\times n}$, is defined as
[…]
Early work on the topic was published in [11, 23] and in several other papers; see [3] for a recent expository paper that provides an overview of the use of the FOV, its history and development, and a comprehensive list of references.
@@
Given a Hermitian positive definite matrix $H\in\mathbb{C}^{n\times n}$, the $H$-norm of a vector $\mathbf{u}\in\mathbb{C}^{n}$ is defined as
$\|\mathbf{u}\|_{H}=(\mathbf{u},H\mathbf{u})^{1/2},$
and the corresponding $H$-norm of a matrix $A\in\mathbb{C}^{n\times n}$ is the induced norm
[…]
If $A$ is nonsingular, its $H$-condition number is defined as
$\kappa_{H}(A)=\|A\|_{H}\|A^{-1}\|_{H}.$
$\begin{bmatrix}F&B^{T}\\ B&0\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \mathbf{p}\end{bmatrix}=\begin{bmatrix}\mathbf{f}\\ \mathbf{g}\end{bmatrix}$, (1)
where $F\in\mathbb{R}^{n\times n}$ is nonsymmetric, $B\in\mathbb{R}^{m\times n}$ has full row rank, and $\mathbf{u},\mathbf{f}\in{\mathbb{R}}^{n},\ \mathbf{p},\mathbf{g}\in{\mathbb{R}}^{m}.$
Significant work has been done on this topic [2, 4, 23, 24, 26], but to the best of our knowledge, the analysis is limited to the situation where 0 is not included in the field of values. Our goal is to perform an FOV analysis for preconditioned saddle-point systems in the case where the origin is included.
Part of our motivation in considering the field of values is that spectral analysis may be limited for this family of linear systems: for nonsymmetric saddle-point systems arising from partial differential equations, the condition number of the eigenvector matrix of the preconditioned matrix typically increases with the matrix dimensions. Let $\mathbf{P}_{j}$ denote all polynomials $p$ of degree $\leq j$ that satisfy $p(0)=1$. Then, considering solving (1) using GMRES in the $H$-norm, if $\mathbf{r}_{k}$ denotes the residual of the $k$th iteration, the inequality
$\frac{\|\mathbf{r}_{j}\|_{H}}{\|\mathbf{r}_{0}\|_{H}}\;\leq\;\kappa_{H}(V)\,\min_{p\in\mathbf{P}_{j}}\max_{k=1,\dots,n+m}\bigl|p(\lambda_{k})\bigr|,$ (2)
where $V$ is the best $H$-conditioned matrix of eigenvectors of the saddle-point matrix and $\{\lambda_{k}\}$ are its $n+m$ eigenvalues, may not capture the possibility of the iteration counts being independent or nearly independent of the matrix dimensions. In the context of the Navier-Stokes equations, for example, this happens even for a large viscosity coefficient (see Remark 4.1). Thus, an analysis of the eigenvalues of the preconditioned matrix is often insufficient to theoretically prove scalability (in situations where it is expected) of an iterative method for such linear systems.
@@
[7] Let $\mathbf{r}_{k}=\mathbf{b}-A\mathbf{x}_{k}$ be the residual of the $k$-th iteration, $\mathbf{x}_{k}$, of GMRES with respect to the $H$-norm applied to the linear system $A\mathbf{x}=\mathbf{b}$, and let $\mathbf{P}_{j}$ denote all polynomials $p$ of degree $\leq j$ that satisfy $p(0)=1$. Then,
$\frac{\|\mathbf{r}_{j}\|_{H}}{\|\mathbf{r}_{0}\|_{H}}\leq\min_{p\in\mathbf{P}_{j}}\|p(A)\|_{H}\leq(1+\sqrt{2})\min_{p\in\mathbf{P}_{j}}\max_{z\in W_{H}(A)}|p(z)|.$
A challenge is that when $0\in W_{H}(A)$, we have $\min_{p\in\mathbf{P}_{j}}\max_{z\in W_{H}(A)}|p(z)|=1$, and Theorem 1.4 fails to provide a useful bound on GMRES convergence in this case. The presence of a zero in the field of values is, in fact, common in saddle-point systems: the (2,2)-block of a saddle-point system preconditioned with a block-diagonal matrix can be 0. Recently, Crouzeix and Greenbaum [7] defined a convex region with a circular hole and showed that it is a spectral set. This can be used to analyze cases when zero is included in the field of values.
In [7] it is shown that if $\Omega_{CG}$ represents a domain constructed as $W_{H}(A)$ with a disk about the origin removed that has radius $1/w$, where $w$ denotes the numerical radius of $A^{-1}$, then $\Omega_{CG}$ is a $(3+2\sqrt{3})$-spectral set for $A$. This bound can be improved to $2+\sqrt{7}$ if a smaller disk about the origin of radius $1/\|A^{-1}\|$ is removed from $W_{H}(A)$; see [7, Theorem 2] and [18], which further refines the discussion of available bounds for a few cases of interest.
@@
[7, 18] Let $A$ be a matrix of the same dimensions as $H$. Then, $\Omega_{CG}=W_{H}(A)\cap\{z\in C:|z|\geq\|A^{-1}\|_{H}^{-1}\}$ is a $(2+\sqrt{7})$-spectral set for $A$.
@@
In this paper, we consider a special family of saddle-point linear systems that arise from discretization of fluid flow problems. We include in our discussion block-diagonal preconditioners and certain block-triangular preconditioners for which no previous FOV analysis is available. On the other hand, our analysis has some limitations compared to the well-studied FOV analysis that excludes the origin. For example, in [24], scaling is effectively used to allow for applying FOV analysis to the discrete Navier-Stokes equations with a small viscosity coefficient when the field of values does not include the origin. In our analysis we are not able to utilize scalings in the same manner, and we require the skew-symmetric part of the linear system to be small norm-wise.
The remainder of this paper is structured as follows. In Section 2 we review some useful results in the literature and present an analysis that deals with zero in the field of values. In Section 3 we specialize our results to saddle-point systems. In Section 4 we discuss a few examples of relevant applications and present some numerical results. Finally, we draw some conclusions in Section 5.
@@ 2.1 Preliminaries
Let us present a few known results that we will use in our analysis. Some of the definitions and results that follow are specialized to real matrices and vectors.
@@
[21, Page 311] For two symmetric positive definite matrices $H_{1}\in\mathbb{R}^{n\times n}$ and $H_{2}\in\mathbb{R}^{m\times m}$, we define the $(H_{1},H_{2})$-norm for a matrix $M\in\mathbb{R}^{m\times n}$ as
[…]
The following equalities, given in [24, Eq. (2.4)], are immediate from Definition 2.1:
@@
The following additional properties from [24] are useful for our analysis.
Lemma 2.2 ([24, Lemma 2.1]).
@@
Let $\mathbf{P}_{n}$ denote the set of polynomials $p$ of degree at most $n$ with $p(0)=1$. For a compact set $S$ in the complex plane, with the origin not included in or surrounded by $S$ and no isolated points, define
$E_{n}(S)=\min_{p\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{P}_{n}}}\max_{z\in S}|p(z)|$
@@
Let $g(z)$ be the Green’s function associated with $S$, defined in the exterior of $S$, satisfying $\nabla^{2}g=0$ outside of $S$, $g(z)\to 0$ as $z\to\partial S$, and $g(z)-\log|z|\to C$ as $|z|\to\infty$ for some constant $C$. Then,
@@ 2.2 Sufficient Conditions
Given constants $a,b,c>0$ with
$bc<1,$ (3)
and a convergence tolerance $\varepsilon>0$, for any nonsingular $n\times n$ matrix $A$ and positive definite $H$ of the same dimensions that satisfy
@@
$\|(HA-A^{T}H)/2\|_{H,H^{-1}}\leq c$ (4c)
for these $a$, $b$, and $c$ values, and any $n$-vector $\mathbf{f}$ and an initial guess $\mathbf{x}_{0}$, there exists some integer $m\geq 1$ that depends on $a$, $b$, and $c$, such that the residual $\mathbf{r}_{m}=\mathbf{f}-A\mathbf{x}_{m}$ generated by $m$ steps of GMRES applied to the linear system $A\mathbf{x}=\mathbf{f}$ in the $H$-norm satisfies
$\|\mathbf{r}_{m}\|_{H}/\|\mathbf{r}_{0}\|_{H}\leq\varepsilon.$
@@
$|Im(z)|$ $\leq\max_{\mathbf{x}\in\mathbb{C}^{n}}\left|\left(\frac{\mathbf{x}^{*}HA\mathbf{x}}{\mathbf{x}^{*}H\mathbf{x}}-\left(\frac{\mathbf{x}^{*}HA\mathbf{x}}{\mathbf{x}^{*}H\mathbf{x}}\right)^{*}\right)\Big/2\right|$
$=\max_{\mathbf{x}\in\mathbb{C}^{n}}\left|\left(\frac{\mathbf{x}^{*}(HA-A^{T}H)\mathbf{x}}{2\mathbf{x}^{*}H\mathbf{x}}\right)\right|$
@@
Notice that the last inequality holds because $HA-A^{T}H$ is skew-symmetric, and hence its field of values lies on the imaginary axis [22, Property 1.2.5].
We then have
@@
$\frac{\|\mathbf{r}_{j}\|_{H}}{\|\mathbf{r}_{0}\|_{H}}$ $\leq\min_{p\in\mathbf{P}_{j}}\|p(A)\|_{H}\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2+\sqrt{7})}\min_{p\in\mathbf{P}_{j}}\max_{z\in\Omega_{CG}}|p(z)|.$
[…]
Figure 1: The shaded region is $\Omega_{D}$ when conditions (3) and (4a)–(4c) of Lemma 2.4 hold
Since Condition (3) holds, the origin is not surrounded by $\Omega_{CG}$, and it follows from Theorem 2.3 that there is always a polynomial (of some degree) with value 1 at the origin that has a maximum magnitude strictly less than 1 on the closure of this set and hence GMRES converges with an asymptotic rate given by $\exp(-g(0))<1$, where $g$ is the Green’s function of this set with a pole at $\infty$ [6, 7]. ∎
@@
If (3) in Lemma 2.4 does not hold, the iterative solver may still converge but we cannot prove convergence using our technique of proof. Specifically, it is immediate to see that $\Omega_{CG}$ is connected and due to the maximum modulus principle, we can only obtain $\min_{p\in\mathbf{P}_{j},p(0)=1}\max_{z\in\Omega_{CG}}|p(z)|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=}1$, which does not indicate convergence; see Figure 2 for a graphical illustration.
[…]
Figure 2: The shaded region is $\Omega_{D}$ when $bc\geq 1$ (i.e., when (3) in Lemma 2.4 is violated)
@@ 2.3 Scope and Limitations of the Analysis
The use of quadratic forms to establish the notion of field of values-based bounds goes back to early work that studied residual-minimizing iterative methods [12, Theorem 3.3]. We provide here the widely used definition of FOV equivalence.
@@
Given two nonsingular matrices $M,A\in\mathbb{R}^{n\times n}$, and a symmetric positive definite matrix $H\in{\mathbb{R}}^{n\times n}$, $M$ is $H$-field-of-values equivalent to $A$ if there exist positive constants $\alpha,\beta$ independent of $n$ such that
$\alpha\leq\frac{(MA\mathbf{x},\mathbf{x})_{H}}{(\mathbf{x},\mathbf{x})_{H}},\ \frac{\|MA\mathbf{x}\|_{H}}{\|\mathbf{x}\|_{H}}\leq\beta,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\qquad\forall\mathbf{x}\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}.}$ (5)
@@
In contrast, in our case, (3) requires $bc$ to be small. While scaling reduces one of $b$ or $c$, it increases the other. Therefore, a simple scaling strategy does not work in the case we are considering, which reveals a limitation of our analysis. We note that condition (4b) is rather standard by norm equivalence considerations (see, for example, [24, Lemma 2.3]). It is condition (4c) that seems to present the difficulty, because it requires the skew-symmetric part of the operator to be smaller than the radius of the inner disk; see Figure 1. Therefore, practically speaking, our analysis is limited to cases where the preconditioned matrix is only mildly nonsymmetric.
However, we note that this lemma may be improved to allow for looser conditions by using a more sophisticated analysis.
@@
The field of values of $A$ is contained in the convex hull of two disks centered at 1-1 with radius $1/4$ and at $2$ with radius $1.2$. Note that while the union of these two disks contains the field of values of $A$ independent of the dimension, the field of values itself does depend (mildly) on $n$. The inverse $A^{-1}$ is available analytically, and it can be shown that $\|A^{-1}\|_{2}^{-1}\to\frac{3}{4}$ as $n\to\infty$; see, for example, [20] for useful relevant results for Toeplitz matrices. For a finite value of $n$, the reciprocal of the norm of $A^{-1}$ needs to be computed numerically, and we have experimentally observed that it is bounded between $0.74$ and $0.76$ for relatively modest values of $n$.
We provide a graphical illustration in Figure 4. Here $c=1.2$ and $b\geq\frac{1}{0.76}$. The condition (3) is violated, but GMRES would still converge for a linear system with the matrix $A$ because $\Omega_{CG}$ does not surround/include the origin. A more careful analysis that tracks the boundary of the FOV (see, e.g., [25]) might result in conditions that are easier to satisfy.
In [18] there is a detailed discussion and a number of examples, including ones of block diagonal matrices similar to the one in Figure 4, where the field of values is divided into two pieces by removing a disk.
@@
While the limitation we have noted is considerable, our analysis substantially broadens the scope of preconditioners for which FOV analysis can be carried out. In particular, in terms of the quantities of Definition 2.5, our analysis makes it possible to consider the scenario that there exists some nonzero $n$-length real vector $\mathbf{x}$ for which
$\frac{(MA\mathbf{x},\mathbf{x})_{H}}{(\mathbf{x},\mathbf{x})_{H}}\leq 0.$ (6)